# Forecasting Trends in Food Security: a Reservoir Computing Approach

### v2

@@ Abstract
Early warning systems are an essential tool for effective humanitarian action. Advance warnings on impending disasters facilitate timely and targeted response which help save lives, livelihoods, and scarce financial resources. In this work we present a new quantitative methodology to forecast levels of food consumption for 60 consecutive days, at the sub-national level, in four countries: Mali, Nigeria, Syria, and Yemen. The methodology is built on publicly available data from the World Food Programme’s integrated global hunger monitoring system which collects, processes, and displays daily updates on key food security metrics, conflict, weather events, and other drivers of food insecurity across 90 countries (https://hungermap.wfp.org/). In this study, we assessed the performance of various models including ARIMA, XGBoost, LSTMs, CNNs, and Reservoir Computing (RC), by comparing their Root Mean Squared Error (RMSE) metrics. This comprehensive analysis spanned classical statistical, machine learning, and deep learning approaches. Our findings highlight Reservoir Computing as a particularly well-suited model in the field of food security given both its notable resistance to over-fitting on limited data samples and its efficient training capabilities. The methodology we introduce establishes the groundwork for a global, data-driven early warning system designed to anticipate and detect food insecurity.
@@ Introduction
Conflict [1, 2], climate extremes, and soaring food, fertilizer and energy prices [3] on the heels of an incomplete recovery from the COVID-19 pandemic [4] have created a food crisis of unprecedented proportions [5]. The war in Ukraine further complicated the situation with millions of people a step away from famine. In such an uncertain environment it is incumbent upon humanitarian agencies such as the World Food Programme to deploy effective early warning systems that monitor food security conditions in the most vulnerable countries [6]. Early warning systems serve as the foundation for preparedness and quick response to potential food crises, whether man-made or natural disasters. They help humanitarian organizations better target assistance to where it is needed the most, hence minimizing duplication and waste. Forecasting systems are an integral part of early warning systems, allowing them to anticipate potential hazards and issue alerts in a timely and effective manner.
@@
Figure 1: (a)-(e): the figure shows the time series of the data used in constructing the forecasting methodology. The target variable, highlighted by the blue dashed curve in (a), is the regional prevalence of insufficient food consumption extracted from the Food Consumption Score (FCS). In addition to the historical values of the target, the methodology incorporates predictors coming from another food security indicator, climate, conflict, and economic data. External datasets with known future values, including crop calendars and Ramadan days, are also considered (e). While the figures are based on data from Mali, the framework remains consistent across all countries. A detailed breakdown of the data for each country can be found in Table 1, with comprehensive information on data sources available in the methods section. (f) The map displays the first administrative level boundaries in the four tested countries, where grey dashed polygons indicate regions that were excluded due to data unavailability.
@@
A robust approach to creating accurate forecasting tools is to leverage modern time series prediction methods that are based on Machine Learning (ML) [7]. These methods can be applied to a growing number of available data streams, and have proven to be successful across a range of diverse fields. Examples of these fields include monitoring epidemics [8, 9, 10, 11, 12, 13], predicting financial markets [14, 15, 16, 17], tracking energy consumption [18] and forecasting weather patterns and climate change impacts [19, 20]. The scientific community has been studying and modeling food security for many decades, and recent research has shown that ML can also be used in this context [21, 22, 23]. By using ML, it is possible to develop complex, data-driven models with minimal feature engineering, thereby avoiding the need for an in-depth understanding of the underlying processes, provided that enough training data is available. This makes ML an attractive candidate for the food security context, where the problem is complex and reliant on local conditions, making it challenging to generalize any claims that may lead to a knowledge-based model.
For instance, the Food and Agriculture Organization of the United Nations (FAO) has investigated long-term forecasts at the country level for several years [24, 25], while nowcasting food security from secondary data has been attempted for households in Uganda [26, 27] and for sub-national units in multiple at-risk countries [22] and comparable work has also been carried out in Malawi [28]. Studies carried out by the World Bank [29, 30] and Westerveld et al. [31] have focused on forecasting transitions between different phases of food security, while a recent study from the French National Centre for Scientific Research (CNRS) has investigated use of ML for the prediction of food consumption[32]. Researchers from UCLA recently used remotely sensed soil moisture data and food prices to predict changes in food security [33]. Even the use of relevant text-based features from news streams has been successfully attempted to forecast food insecurity as defined by the Integrated Phase Classification (IPC) up to 12 months ahead [23]. Notably in [34] authors used a panel vector-auto regression to model and forecast IPC 1, 2 and 3+ phases in 15 countries, and then applied a LASSO to identify the most important drivers. Finally in [21] authors used XGBoost to build a 30-day forecaster on a similar data set up used in this work.
The World Food Programme regularly collects data on food security in countries where it operates [35]. The methodology we present in this manuscript builds upon data sourced from the Real-Time Monitoring (RTM) system, developed by the organization over the last decade and publicly available via HungerMapLIVE [36]. Collected through Computer Assisted Telephone Interviews (CATI)[37], the data provides daily updates on food security indicators at the household level in near 40 countries. One such indicator is the Food Consumption Score (FCS), a composite measure that assesses the frequency of consumption of different food groups in the seven days prior to the survey, categorizing households as having poor, borderline, or acceptable food consumption. Households with insufficient food consumption refers to those with poor or borderline food consumption and is the target variable for our study. In contrast to other food security monitoring initiatives like IPC, CH, or FEWSNet, WFP’s RTM program stands out due to several significant distinctions. Other programs release country updates at intervals of six months or more while the RTM program provides daily updates throughout the entire year. This unique feature enables real-time monitoring of the dynamic changes in food security levels, setting it apart as a more responsive and timely tool for assessing and addressing evolving challenges. Furthermore, the underlying process to generate the indicators in the RTM is purely quantitative, and does not rely on domain expert consensus or qualitative observations.
The main goal of this research is to develop a forecasting methodology capable of predicting insufficient food consumption, as presented in WFP’s RTM system, over a span of 60 consecutive days. This has been attempted in four countries, Mali, Nigeria, Syria, and Yemen, at the sub-national level, using secondary data that represent the key-drivers of food insecurity: conflict, extreme weather events and economic shocks [6]. A successful forecasting methodology developed using the RTM data, building on the unique properties of the monitoring system, would therefore be complementary to the reported approaches, and prove valuable for quantitative timely insights at any given point in time. This includes anticipating levels of food insecurity following natural disasters, political or economic shocks, and while the lean season is unfolding when other information may be scarce. The versatility and flexibility of such methodology would ensure its utility all year around, and would not be limited to the predefined release dates of monitoring reports. Moreover, in other methodologies that extend their forecasts over longer periods, the longer duration is typically a result of the increased time granularity of the data. In such cases, a 12-month forecast corresponds to predicting only one or two data points into the future. Our methodology is designed to replicate the complete 60-day time series of insufficient food consumption, thereby estimating the curve’s behavior throughout the entire time window and going beyond offering a single numerical output at the conclusion of the period. In order to increase the transparency and explainabilty of the model, an aspect that has been identified as crucial in the literature [38, 39, 40], we further implement a simple method for dynamic feature selection. This allows us to assess the importance of different variables in the model and their impact on the prediction.
We compare a broad range of approaches, covering both classical statistical methods and state-of-the-art deep learning techniques. The model we identify to be the most suitable for the problem of forecasting trends of food insecurity is based on a Reservoir Computing (RC)[41, 42, 43], a type of recurrent neural network (RNN), where only the last readout layer is trained. This approach, recently attracting considerable attention and application across various domains, offers a fast and efficient training procedure while mitigating challenges such as the vanishing gradient problem. Despite its simplicity RC has proven to be a valid alternative to advanced machine learning prediction techniques particularly in data-scarce contexts as can be found in [44, 45, 46] when applied to synthetic data. Specifically our methodology relies on an ensemble of RC models, exploiting the low resource requirements of training RC and stabilizing predictions. This was also recently attempted in [47] where RC is shown to outperform LSTMs on forecasting food prices. To assess the goodness of our approach, we conducted a comprehensive evaluation by comparing the Root Mean Squared Error (RMSE) of various methodologies: ARIMA, CNNs and LSTMs. Additionally, we referred to the results presented in [21], where the authors developed an XGBoost forecaster. The findings presented in this manuscript highlight Reservoir Computing (RC) as a robust option for forecasting applications in the field of food security. Notably, RC outperformed the other tested algorithms, particularly in situations of more dramatic deterioration.
@@
Figure 2: 60-day forecasts were generated using the RC, CNN, LSTM, and ARIMA models for four specific sub-national regions: Yemen, Syria, Mali, and Nigeria. In the visual representation, the blue curve represents the actual data, while each of the other curves depicts the prediction of one of the models. These predictions cover the time period starting on the first of May 2023, with each model trained on all available data up to that point.
@@ Results
Figure 3: (a)-(c) Performances of the LSTM, CNN and ARIMA models measured in median RMSE, aggregated in three different ways. (a) country (b) prediction lag (c) behavior of the target time series. In (c) the inset plot shows a comparative analysis of each model’s performance against the RC model based on the same metric, and the distribution of the samples over the different types of behaviour. (d) Average time cost for training each of the models used. The time was recorded during creation of the forecasts used in the main analysis. The RC time corresponds to the training of 100 separate models, as done to produce the forecasts and measure the performances e Comparison of the RC model’s performance to the XGBoost model [21] on five 30-day time windows between October 2021 and February 2022 as shared by the authors.
@@
The target variable of our modelling is the time series of the sub-national prevalence of households with insufficient food consumption. An example is shown in the left panel of figure 1, together with time series examples of a sub sample of the secondary data. Our choice is inspired by the approach used in previous works [21, 22] and makes use of available datasets describing conflict, economic shocks, and weather events. More specifically these consist of rainfall and vegetation levels, spikes in market food prices, food and headline inflation, currency exchange rates and number of fatalities due to different kinds of conflict. A detailed description of the input data is provided in the Methods section. Since Islam is the main religion in all four analysis countries, the holy month of Ramadan (fasting) is expected to significantly influence insufficient food consumption (we can see this in figure 5 in the Methods section). Therefore, the dates for Ramadan are used as an external input in the RC model, since it is available not only for the past, but also for the prediction period. To standardize all input variables to the same format of sub-national daily time series we performed some basic pre-processing steps. These include linear interpolation of variables which are only available on monthly basis, aggregating information to the desired sub-national level, averaging intermittent variables over a rolling window as well as de-noising and smoothing all data. Given the stochastic nature of the RC algorithm, to generate our output we repeat the prediction procedure 100 times with different seeds, with the median of all results then considered the final prediction.
@@
A comparison between the predicted curves obtained with the different tested methodologies and the actual data is shown in the plots in figure 2 where the blue curves are the actual data representing the sub-national curves of insufficient food consumption, the colored curves representing the forecasts. The objective of this work is to test the goodness of the predictions for 60 consecutive days. To ensure a comprehensive comparison among all considered algorithms and to assess their robustness across all seasonal trends, we have evaluated their performances in 12 60-day splits. The forecasts for each split starter at the beginning of every month, covering the period from June 2022 to May 2023. To simulate the use of such methodology in a production environment we employed a walk-forward validation scheme, meaning the model was retrained for every new forecast adding the data of the previous month, therefore expanding the training moving forward in time. At this stage retroactive delays in data updates were not factored in and the models were trained on all available data at present (refer to Supplementary Material for additional details). For every split the hyper-parameters of the different algorithms were updated and selected through an extensive grid search of the performances on previous splits. Furthermore the algorithms were allowed to choose the subset of input variables that were found to be more informative, i.e. minimized the errors similarly to what was measured in [22]. This approach represents a first step towards adding to the forecasting methodology a method to explain the role of the input variables and their importance in generating the forecasts. In figure 4 of the Methods section the subsets of input variables are reported along with the frequency of their selection. It is clear that most of the predictive power comes from the auto-regressive element, the rCSI indicator and the seasonal external variables. A detailed description of the training procedure is found in the Methods section.
The aggregate errors are shown in figure 3 and are the result of ca. 70.000 CPU hours of calculations accumulated over several 10.000 runs for each country and model on the German Aerospace Agency’s High Performance Cluster. This extensive computation period was essential to thoroughly assess the performances achieved with a vast array of hyper-parameter combinations across all tested splits and countries, ensuring a fair comparison between methodologies. It’s noteworthy that the actual training times, on a local standard machine, for an individual hyper-parameter combination, illustrated in Figure 3(d), typically fall under 1 minute. As illustrated in the figure, our analysis of the performances includes a comprehensive study from multiple angles, emphasizing systematic differences in the outputs. To achieve this, we aggregated errors by country, by classifying the single splits of the underlying data according to their behavior, and explored the temporal evolution of errors. Testing the methodologies on different countries is a way to assess the algorithms’ robustness to different sets of available input data sets as well as diverse underlying relationships with the target variable. A complete reference of the available input data sets per country is found in table 1.In figure 3a we can see how RC clearly outperforms the two deep learning algorithms LSTMs and CNNs in all countries, and is either better or comparable to ARIMA. Perhaps surprisingly the simple ARIMA algorithm outperforms the two more sophisticated LSTMs and CNNs. This type of results is in line with the findings in other studies [48, 49, 50, 51, 47] where ARIMA outperforms or yields comparable results to more advanced algorithms, and is likely to depend on the high amount of noise present in the data as well as its scarcity. The temporal evolution of the error is shown in Figure 3b: initially CNNs and ARIMA exhibit strong performance in the early time steps of the predictions, however over time RC distinctly outperforms all benchmarks, showcasing its increasing advantage as the prediction horizon extends. Additionally, Figure 3c illustrates the third and arguably the most crucial type of aggregation. The figure demonstrates a widening difference in errors among the models as the target variable exhibits more extreme variations, with the RC consistently outperforming all benchmarks by larger margins. The reason for the increase in errors across all models for larger variations in the data is attributed to the rarity of such behavior. The training dataset is indeed unbalanced and skewed toward stable curves as shown in the inset of figure 3c. Remarkably RC’s performance, although in relative terms, experiences a substantial increase. This property is exceptionally valuable, and optimizing the methodology for these less common curve types will be the focus of future work. In figure 3d, we observe the training time for all tested algorithms on a single split, using the hyperparameters selected from the grid search. The comparison reveals that, despite employing an ensemble of 100 models for Reservoir Computing (RC), its training time remains comparable to ARIMA and is significantly more efficient than LSTMs. While the CNN achieves the shortest training time, this advantage is counterbalanced by a considerably inferior performance. The relatively long training times obtained using ARIMA depend on the necessity to train a distinct ARIMA model for each region in a country. In contrast, other methodologies enable simultaneous training, allowing for concurrent forecasting across all regions.
To further evaluate our proposed approach, we conducted a benchmark on the same dataset employed and shared in in [21]. In their study, the authors implemented an XGBoost forecaster for a consecutive 30-day period, targeting the same variable as addressed in our manuscript. It’s worth noting that XGBoost inherently lacks built-in support for multi-output regression, as its primary design focuses on single-output tasks such as regression and classification. Consequently, to achieve a 30-day prediction horizon, the authors were forced to train and evaluate thirty distinct models. In figure 3e we can clearly see how the RC outperformed the XGBoost in all tested countries. Given XGBoost’s reliance on multiple models, rendering its use in a production environment challenging in terms of maintenance and updates, and the clear results obtained in comparing the performances on a horizon of 30 days, extending the approach to a $60$-day prediction horizon was considered impractical and unlikely to provide substantial additional value.

### v3

# Forecasting trends in food security with real time data

@@ Abstract
Early warning systems are an essential tool for effective humanitarian action. Advance warnings on impending disasters facilitate timely and targeted response which help save lives and livelihoods. In this work we present a quantitative methodology to forecast levels of food consumption for 60 consecutive days, at the sub-national level, in four countries: Mali, Nigeria, Syria, and Yemen. The methodology is built on publicly available data from the World Food Programme’s global hunger monitoring system which collects, processes, and displays daily updates on key food security metrics, conflict, weather events, and other drivers of food insecurity. In this study we assessed the performance of various models including Autoregressive Integrated Moving Average (ARIMA), Extreme Gradient Boosting (XGBoost), Long Short Term Memory (LSTM) Network, Convolutional Neural Network (CNN), and Reservoir Computing (RC), by comparing their Root Mean Squared Error (RMSE) metrics. Our findings highlight Reservoir Computing as a particularly well-suited model in the field of food security given both its notable resistance to over-fitting on limited data samples and its efficient training capabilities. The methodology we introduce establishes the groundwork for a global, data-driven early warning system designed to anticipate and detect food insecurity.
@@ Introduction
Conflict [1, 2], climate extremes, and soaring food, fertilizer and energy prices [3] on the heels of an incomplete recovery from the COVID-19 pandemic [4] have created a food crisis of unprecedented proportions [5]. The war in Ukraine further complicated the situation with millions of people a step away from starvation. In such an uncertain environment it is incumbent upon humanitarian agencies to deploy effective early warning systems that monitor food security conditions in the most vulnerable countries [6].The World Food Programme (WFP) operates in emergency contexts and its operations include direct delivery of food assistance, cash based transfers, nutrition support, and are characterized by rapid response. This is done by leveraging an extensive logistic network in partnership with other UN agencies, NGOs, government counter parts and local communities, and requires swift assessments of local needs and vulnerabilities. In this context early warning systems serve as the foundation for preparedness and quick response to potential food crises, whether man-made or natural disasters. They help humanitarian organizations better target assistance to where it is needed the most, hence minimizing duplication and waste. Forecasting systems are an integral part of early warning systems, allowing them to anticipate potential hazards and issue alerts in a timely and effective manner. A robust approach to creating accurate forecasting tools is to leverage modern time series prediction methods that are based on Machine Learning (ML) [7]. These methods can be applied to a growing number of available data streams, and have proven to be successful across a range of diverse fields. Examples of these fields include monitoring epidemics [8, 9, 10, 11, 12, 13], predicting financial markets [14, 15, 16, 17], tracking energy consumption [18] and forecasting weather patterns and climate change impacts [19, 20].
@@
Figure 1: Input Data: the figure shows the time series of the data used in constructing the forecasting methodology. The target variable, highlighted by the blue dashed curve in (a), is the regional prevalence of insufficient food consumption extracted from the Food Consumption Score (FCS). In addition to the historical values of the target, the methodology incorporates predictors coming from another food security indicator, (b) climate , (c) conflict, and (e) economic data. External datasets with known future values, including crop calendars and Ramadan days, are also considered (d). While the figures are based on data from Mali, the framework remains consistent across all countries. A detailed breakdown of the data for each country can be found in Table 1, with comprehensive information on data sources available in the methods section. (f) The map displays the first administrative level boundaries in the four tested countries, where grey dashed polygons indicate regions that were excluded due to data unavailability.
@@
The scientific community has been studying and modeling food security for many decades and recent research has shown that ML can also be used in this context. By using ML, it is possible to develop complex, data-driven models with minimal feature engineering, thereby avoiding the need for an in-depth understanding of the underlying processes, provided that enough training data is available. This makes ML an attractive candidate for the study of food security, where the problem is complex and reliant on local conditions, making it challenging to generalize any claims that may lead to a knowledge-based model. An ML approach, if well designed, is able to reconstruct the underlying causal pathways without pre-established modelling choices, offering an alternative to the modelling from first principles. There exists a comprehensive body of literature on methods that leverage ML techniques, either standalone or in combination with other approaches.
For instance, the Food and Agriculture Organization of the United Nations (FAO) has investigated long-term forecasts at the country level for several years [21, 22]. In [23, 24] ML algorithms were used to classify households based on their caloric intake in Uganda, from data collected by the Bureau of Statistics. Studies carried out by the World Bank focus on the modeling and forecasting of Integrated Food Security Phase Classification (IPC) phases based on data from the Famine Early Warning Systems Network (FEWS.NET) across fifteen countries [25, 26, 27]. The studies include attempts to identify the onset of food crises by building a binary crises indicator from the IPC phases and reconstruct the national share of population living in districts labeled as IPC3+[25]. Moreover in [26] forecasts sub-national IPC phases 1, 2, and 3+ employing a panel vector-auto regression model and using the Least Absolute Shrinkage and Selection Operator (LASSO) technique to pinpoint the most critical drivers of food security. Notably in [28], based on the same FEWS.NET data, the authors extract significant text-based features from news articles accessed via Factiva. In their work they demonstrate that the text features greatly enhance the predictive capabilities of a Random Forest model, thus exploiting the potential of an abundant data stream that has not been thoroughly explored in the context of food insecurity.
@@
Figure 2: Forecasts and Data: (a) 60-day forecasts examples generated using the RC, CNN, LSTM, and ARIMA models for four specific sub-national regions: Yemen, Syria, Mali, and Nigeria. In the visual representation, the blue curve represents the actual data, while each of the other curves depicts the prediction of one of the models. (b) The observed distribution of the variation of the prevalence of insufficient food consumption on the 60-day windows used to train and test the algorithms. We can appreciate how the dataset is biased towards curves of small variation.
Using data from the Living Standards Measurements Study (LSMS) in [29], a LASSO regression was used to forecast food security indicators such as the Food Consumption Score (FCS), the Household Dietary Diversity Score (HDDS) and the reduced Coping Strategy Index (rCSI), in Malawi. In [30] an XGBoost algorithm is trained on secondary data in Ethiopia spanning a wide range of available data sets to forecast the future behaviour of the FEWS.NET IPC phases, classified as deteriorations, improvements and no change. A recent study from the French National Centre for Scientific Research (CNRS) has investigated use of Deep Learning approaches to forecast FCS and HDDS indicators as collected annually from the government in Burkina Faso. While researchers from UCLA recently used remotely sensed soil moisture data and food prices to predict changes in food security, using data provided by IPC in 7 countries [28]. Leveraging real-time data collected by the World Food Programme, in [31] authors developed a cross-country prediction model to Nowcast FCS and rCSI indicators across sub-national units in multiple countries, while in [32] authors trained 30 distinct XGBoost models to construct a 30-day forecasting model for FCS. In this manuscript we used the same target and secondary variables used in [31, 32]
The World Food Programme regularly collects data on food security in countries where it operates. The methodology we present in this manuscript builds upon data sourced from the Real-Time Monitoring (RTM) system, developed by the organization over the last decade and publicly available via HungerMapLIVE [33]. Collected through Computer Assisted Telephone Interviews (CATI)[34], the data provides daily updates on food security indicators at the household level in near 20 countries. One such indicator is the Fpood Consumption Score (FCS), a composite measure that assesses the frequency of consumption of different food groups in the seven days prior to the survey, categorizing households as having poor, borderline, or acceptable food consumption. Households with insufficient food consumption refers to those with poor or borderline food consumption. The aggregation of this variable to the first administrative level in in countries is the target variable for our study. In contrast to other food security monitoring initiatives like IPC, CH, or FEWSNet, WFP’s RTM program stands out due to several distinctions. Other programs release country updates at intervals of four or more months while the RTM program provides daily updates throughout the entire year. This unique feature enables real-time monitoring of the dynamic changes in food security levels, setting it apart as a more responsive and timely tool for assessing and addressing evolving challenges. Furthermore, the underlying process to generate the indicators in the RTM is purely quantitative, and does not rely on domain expert consensus or qualitative observations.
The primary aim of this research is to develop a forecasting methodology that can predict insufficient food consumption, as described in the WFP’s RTM system, over a continuous 60-day period. While national development plans can depend on low-frequency data released on fixed dates, the strategic allocation and distribution of resources in emergency scenarios require more frequent updates, such as monthly or weekly, and constant prioritization for optimal impact. Our early warning system is specifically designed for the dynamic and rapidly changing conditions of humanitarian aid operations in crisis settings. The forecasts are updated daily, enhancing their versatility and flexibility, making them continuously useful throughout the year, unlike other methodologies that are restricted by fixed report release schedules. This feature supports frequent re-targeting and optimization, which are crucial for effective humanitarian responses. Given the logistical challenges in hardship conditions where resource distribution can take up to two weeks, our 60-day forecasts are valuable for assessing the risk of deterioration and enabling planning based on updated priorities. Therefore, a 60-day forecasting horizon, with frequent updates, is considered operationally valuable for managing resource distribution programs effectively.
Figure 3: Forecasting Aggregated Performances: Performances of the LSTM, CNN and ARIMA models measured in median RMSE, aggregated in three different ways: Foretasting time step a, variation of the target variable b, per country c. Figure a shows how the RC model outperforms the other methodologies after the 15th forested step with an aggregated RMSE at the end of the 60-day window of 4.9 percentage points. In figure b we see RC tends is consistently the best performer when aggregating on curves according to the variation of their target variable $\Delta=\textbf{fcs}_{60}-\textbf{fcs}_{0}$. This is more evident for high values of $\Delta$, that indicate a sharp increase of levels of insufficient food consumption. The barplot in figure c aggregates the error per country. Despite the fact the RC is among the top performers for every country there is not a clear preferred methodology. This is due to the fact that regional time series belonging to the same country can have very different behaviours and have no clear national characteristic.
We tested out the methodology in four countries, Mali, Nigeria (north east), Syria, and Yemen, at the first sub-national level, using secondary data that represent the key-drivers of food insecurity: conflict, extreme weather events and economic shocks [6]. The selection of dataset categories for secondary variables in our study draws inspiration from the choices made in previous research [26, 30, 25, 27, 28, 29, 31, 32], with some adjustments based on availability and access convenience. A detailed description is found table 1. Our forecasting methodology innovates by leveraging RTM data as the primary target, capitalizing on the unique aspects of the monitoring and incorporating several desirable features that are particularly beneficial for informing humanitarian assistance. In order to increase the transparency and explainability of the model, an aspect that has been identified as crucial in the literature [35, 36, 37], we further implement a simple method for dynamic feature selection. This allows us to assess the importance of different variables in the model and their impact on the prediction.
The model we identify as the most suitable for forecasting trends in food insecurity is based on Reservoir Computing (RC) [38, 39, 40], a type of recurrent neural network (RNN). Several features make this approach particularly appealing in this context. First, when training the RC, only the parameters of the final layer are "learned," while all other layers are kept constant. This simplification effectively makes RC equivalent to a linear regression, eliminating the need for gradient descent and reducing the number of parameters to learn drastically, speeding up the process and making it suitable for data-constrained contexts. Despite this simplification, the untrained layers and activation functions maintain the model’s ability to capture non linear and complex patterns in the data. Additionally, RC allows for the dynamic treatment of secondary variables which are iteratively estimated along with the target variable. The model learns to project these secondary variables into the future, facilitating a straightforward treatment of multidimensional time series. Furthermore, the architecture of RC enables the easy addition of exogenous variables whose future values are known and do not need to be estimated. This feature enhances the model’s flexibility and accuracy by incorporating external information directly into the forecasting process. Despite its simplicity RC has proven to be a valid alternative to advanced machine learning prediction techniques particularly in data-scarce contexts as can be found in [41, 42, 43] when applied to synthetic data. Specifically our methodology relies on an ensemble of RC models, exploiting the low resource requirements of training RC and stabilizing predictions. This was also recently attempted in[44] where RC is shown to outperform LSTMs on forecasting food prices.
To assess the goodness of our approach, we conducted a comprehensive evaluation by comparing the Root Mean Squared Error (RMSE) of various methodologies, that span deep learning, machine learning and classical statistical model: ARIMA, CNN, LSTM and the XGBoost from a previous attempt [32]. Inspired by the approach in [30], we also developed a simple classifier to label future behavior as Deteriorating, No Change, or Improving based on the forecasted values. Our results demonstrate that the RC model outperforms other tested approaches in both types of tasks, and in particular in anticipating dramatic deterioration. This is notable given that curves that show deterioration and severe deterioration are less common in the dataset than the stable curves. For these reasons we believe that the findings presented in this manuscript highlight Reservoir Computing (RC) as a robust option for forecasting applications in the field of food security, where high frequency data is scarce and skewed towards stable behaviour, and for producing early warning signals of impending food crises.
Figure 4: Classification Performance: This figures compares the performances of the ARIMA, CNN, LSTM and RC models on their ability to distinguish different classes of behaviour: No Change, Improvement and Deterioration as defined in (a). The same criterion in applied to actual curves and predicted curves to label the behaviour. (b) Performances metrics on the classification task. Accuracy on singles classes are computed treating the problem as a binary classification (c) Confusion Matrices of all tested methodologies.
@@ Results
The target variable of our modelling is the time series of the prevalence of households with insufficient food consumption at the first sub-national level. An example is shown in the left panel of figure 1, together with time series examples of a sub sample of the secondary data. Our choice is inspired by the approach used in previous works [32, 31] and makes use of available datasets describing conflict, economic shocks, and weather events. More specifically these consist of rainfall and vegetation levels, spikes in market food prices, food and headline inflation, currency exchange rates and number of fatalities due to different kinds of conflict. A detailed description of the input data is provided in the Methods section. Since Islam is the main religion in all four analysis countries, the holy month of Ramadan (fasting) is expected to significantly influence insufficient food consumption (see the Methods section). Therefore, the dates for Ramadan are used as an external input in the RC model, since it is available not only for the past, but also for the prediction period.
To standardize all input variables to the same format of sub-national daily time series we performed some basic pre-processing steps. These include linear interpolation of variables which are only available on monthly basis, aggregating information to the desired sub-national level, averaging intermittent variables over a rolling window as well as de-noising and smoothing all data. Given the stochastic nature of the RC algorithm, to generate our output we repeat the prediction procedure 100 times with different seeds, with the median of all results then considered the final prediction.
A comparison between the predicted curves obtained with the different tested methodologies and the actual data is shown in the plots in figure 2(a) where the blue curves are the actual data representing the sub-national curves of insufficient food consumption, the colored curves representing the forecasts.In 2(b) we can see how the variation of the curves is distributed in the dataset, where $\Delta=\text{fcs}_{60}-\text{fcs}_{0}$, expressed in percentage points, represents the difference between the last and first values in the forecasting window. A positive $\Delta$ implies that the last value of the window is higher than the first, therefore indicating an increase of food insecurity. On the other hand a negative value indicates an improvement. We refer to curves with positive $\Delta$ as deteriorating curves, while those with a negative $\Delta$ as improving curves. We can see from the figure that the dataset is skewed towards small values of $\Delta$, indicating a majority of stable curves.
Figure 5: Feature Selection: Feature groupings that were considered in the grid search for the RC, CNN and LSTM models to implement a dynamic feature selection.(a) Schema of how the features were classified into 5 different feature groupings. (b-d) frequency with which each feature grouping was selected for in (a) the RC model, (b) the LSTM model and (c) the CNN model.
@@
The objective of this work is to test the goodness of the predictions for 60 consecutive days. To ensure a comprehensive comparison among all considered algorithms and to assess their robustness across all seasonal trends, we have evaluated their performances in 12 60-day splits. The forecasts for each split starts at the beginning of every month, covering the period from June 2022 to May 2023. To simulate the use of such methodology in a production environment we employed a walk-forward validation scheme, meaning the model was retrained for every new forecast adding the data of the previous month, therefore expanding the training moving forward in time. At this stage retroactive delays in data updates were not factored in and the models were trained on all available data at present. In a production environment one must take into consideration the update frequency of the secondary variables as well as the target, that may not happen regularly. In section 2 of the supplementary material we further address this issue. For every split the hyper-parameters of the different algorithms were updated and selected through a grid search of the performances on previous splits. Furthermore the algorithms were allowed to choose the subset of input variables that were found to be more informative, i.e. minimized the errors similarly to what was measured in [31].
The aggregate errors are shown in figure 3 and 4 and are the result of ca. 70.000 CPU hours of calculations accumulated over several 10.000 runs for each country and model on the German Aerospace Agency’s High Performance Cluster. This extensive computation period was essential to thoroughly assess the performances achieved with a vast array of hyper-parameter combinations across all tested splits and countries, ensuring a fair comparison between methodologies. It’s noteworthy that the actual training times, on a local standard machine, for an individual hyper-parameter combination typically fall under one minute for all methodologies (See the supplementary material section 5 for more details). As illustrated in figure 3, our analysis of the performances includes a comprehensive study from multiple angles, emphasizing systematic differences in the outputs.
In figure 3a we start by exploring the temporal evolution of the errors, showing the aggregate error for each forecasting step. We can see how the RC model’s error rate increases at a slower pace compared to all other considered algorithms. While the RC model is initially outpaced up to the 15th forecasting step, it distinctly becomes the preferable option beyond this point. Figure 3b presents the second type of aggregation for the errors: the degree of change in the data. In the figure we observe that the difference between the performances of the RC and the other tested model widens as the target variable undergoes more substantial change, with RC model outperforming all benchmarks by a clear margin under severely deteriorating conditions. We interpret the general increase across all models in errors for larger data variations as due to the infrequency of such extreme changes, as highlighted in figure2 b. Enhancing our approach to better capture these less common types of curves will be the aim of our future work. Figure 3c, shows the results of aggregating the methodologies on different countries, to assess the algorithms’ robustness to different sets of available input data sets as well as diverse underlying relationships with the target variable. A complete reference of the available input data sets per country is found in table 1. 3c reveals that both the RC and the ARIMA models emerge as performing better than other methodologies overall, but no clear preferred choice emerges from this aggregation. We believe that this is due to the fact that regional time series of insufficient food consumption do not show any evident national property or characteristics and can assume very different behaviour, even within successive splits. The good performance shown by the simple ARIMA model, frequently among the top choices in all aggregation, aligns with findings from several other studies [45, 46, 47, 48, 44], where the ARIMA model often surpasses or matches the effectiveness of more sophisticated algorithms. As a further evaluation of our proposed approach we compared the performance on 30-day forecasts with those shown in [32] on the same dataset. The comparison is found in section 6 of the Supplementary Material where we can see how RC once again stands out the best choice.
Figure 6: Ramadan Effect: Prevalence of people with insufficient food consumption averaged over all available regions of Yemen in blue. Time periods of Ramadan shown by areas shaded red.
Inspired by [30], we classified curves into three categories based on their variation: "No Change" for curves with a variation $|\Delta|<0.04$, "Improvement" for a $\Delta<-0.04$, and "Deterioration" for $\Delta>0.04$. Once again $\Delta=\text{fcs}_{60}-\text{fcs}_{0}$ is the measure of change seen in the curves. We also applied the same classification criterion to the forecasted curves, turning the regression problem into a classification. From figure 4 we can see that the RC model shows superior performance across all metrics with the exception of the accuracy on stable curves. Other methodologies tend to label the majority of curves as being stable therefore falling in the accuracy paradox, where high performances depend on the class unbalance of the dataset rather than on correct prediction. When looking at the minority class Deterioration, RC clearly stands out as the best choice in accuracy, recall and precision. It is worth underlying that all models, including the RC, show a conservative behaviour and tend to forecast a stable outcome. We can see this from the very low levels of the recall on the Deterioration class. This is in line with what was seen in [25] where the False positive rates shown by the methodology are one order of magnitude lower than the False negative rates, and is a limitation of the current methodology that will be addressed in future work.
Finally we have studied what subset of secondary variables produced the lowest error for each split that was tested. In figure 5 the subsets of input variables are reported along with the frequency of their selection for every methodology. It is clear that most of the predictive power for all tested methodologies comes from the auto-regressive element, the rCSI indicator and the seasonal external variables Understanding how to optimise the triangulation of the data sources will be the object of future work. We further analyse this important point in the Training Procedure sub-section of the Methods section.